# Defect versus defect: stationary states of single file marching in periodic landscapes with road blocks

### v2

@@ Abstract
Totally asymmetric simple exclusion process (TASEP) sets the paradigm for one-dimensional driven single file motion. We study a periodic TASEP with two “road blocks” or defects of different kinds, one point and another extended, across which particle flows are inhibited. We show how the interplay between particle number conservation and competition between the defects lead to inhomogeneous steady states with localised domain walls (LDW) or static density shocks, whose locations jump discontinuously, indicating a first order transition between these LDW states, as the system passes from being controlled by one defect to the other. When the defects are “competing”, instead of an LDW a pair of delocalised domain walls are found, none of which can penetrate the extended defect. A minimum current principle can be used to identify the dominant defect that controls the domain wall formations.
@@
Single file motion (SFM) implies particle motion along quasi one-dimensional (1D) narrow channels where the particles cannot cross each other due to hardcore repulsion. This was originally introduced by Hodgkin and Keynes hodg to describe ion transport in biological channels. Unidirectional or driven SFM have received increasing attention in the recent past for their wide-ranging interdisciplinary applications. We are interested in the role of bottlenecks and their interplay and competition in controlling the stationary states of driven SFM, a question of potential relevance in wide-ranging systems, e.g., traffic flow  ped1 ; ped2 ; ped3 and colloidal particles flowing through geometric constraints coll1 ; coll2 ; coll3 .
Totally asymmetric simple exclusion process (TASEP) serves as a paradigmatic example of 1D driven SFM. Consisting of 1D lattice with $L$ sites, each site of a TASEP can accommodate at most one particle that can hop unidirectionally if the neighbouring site is empty. It was originally proposed to describe protein synthesis by ribosome translocation along messenger RNA strands in eukaryotic cells mac . Subsequently, it was re-investigated as a simple model for nonequilibrium phase transition in 1D open systems krug ; tom .
@@
In this Letter, we study a particle number conserving TASEP in a ring with one point (PD) and one extended (ED) defects as a conceptual model for competition among road blocks in driven SFM having a fixed number of mobile agents. We elucidate how the stationary density profiles of the TASEP emerge as an interplay between the dominant defect and number conservation. Our principal results are (i) at a qualitative level, the defects lead to domain walls (DW) in the stationary density profiles connecting segments lower and higher than 1/2,when the system is neither nearly empty or nearly full, but rather has a moderate density. This defines a DW phase. In a striking display of defect competition, the DW location jumps discontinuously as the system undergoes a first order transition between PD and ED controlled states. (ii) In the DW phase, generally there is a single localised DW (LDW) for both PD or ED. Nonetheless, the LDW profiles for a PD and ED are fundamentally different from each other. (iii) In the event, the two defects are equi-dominant (in the sense explained below), instead of a single LDW, a pair of delocalised domain walls (DDW) emerges, forming outside ED. (iv) For a nearly empty or filled system, the density everywhere is less or more than 1/2, reminiscent of the low density (LD) or high density (HD) phases of a TASEP with open boundary conditions krug ; tom ; rev1 .
Our results on DWs can be reconciled within a minimum current principle. This study should be useful to understand the phenomenologies of ribosome translocations along closed mRNA loops (circular translation of polysomal mRNA) with clusters of slow codons along which ribosome translocations are inhibited mrna1 ; mrna11 ; mrna2 ; mrna3 ; mrna4 ; mrna5 .
@@
The phases in the model and the associated transitions are controlled by $p,q,n$. The corresponding mean-field phase diagram with $\epsilon_{1}=\epsilon_{2}=1/3$ is shown in Fig. 2, giving the phases in the three segments - $T_{1}$ between $i=1$ to $\epsilon_{1}L$, $T_{2}$ between $\epsilon_{1}L$ to $(\epsilon_{1}+\epsilon_{2})L$ and $T_{3}$ between $(\epsilon_{1}+\epsilon_{2})L$ to $L$, connected serially.
Figure 2: Mean-field phase diagram of the model in the $p-q-n$ space showing the phases in the model. HD and DW (with one LDW) phases are marked. LD phase that lies below the DW phase in the figure is not visible. Points on the gray curved plane lying as a divider of the DW phase have $J_{\text{PD}}=J_{\text{ED}}$ satisfied, where $J_{\text{PD}},\,J_{\text{ED}}$ are the steady state currents in the PD and ED dominated situations respectively (see text below) and have a pair of DDWs fully or partially covering $T_{1}$ and $T_{3}$. In the DW phase region, on the left of this curved surface (i.e., with $J_{\text{PD}}<J_{\text{ED}}$, there is an LDW due to the point defect in either $T_{1},T_{2}$ or $T_{3}$; on the right of this plane (i.e., with $J_{\text{ED}}<J_{\text{PD}}$, one has an LDW due to the extended defect in $T_{1}$ or $T_{3}$ and correspondingly an MC phase in $T_{2}$. In the HD (LD) phase region, one has HD-HD-HD (LD-LD-LD) phase in $T_{1},T_{2},T_{3}$ (see text).
@@
The emergence of DWs, as revealed in our MCS results and observed in the movies sm can happen in two distinct ways. Physically, from the LD-LD-LD phase upon addition of particles, $n$ rises and it eventually reaches a lower threshold, a macroscopically nonuniform steady state in the shape of a DW is formed. A DW, essentially a pile up of particles due to a bottleneck or a defect, should form behind the defect, point or extended, as the case may be, and then start to grow as $n$ rises further. While this remains true independent of the defect, our MCS results reveal a striking aspect - a DW that forms behind a PD can be anywhere in the TASEP ring covering all of $T_{1},\,T_{2},\,T_{3}$, being controlled by $n$, whereas a DW, which forms behind an ED, never enters $T_{2}$! This leads to complex density profiles when the two defects are “competing”; see below and also Movie 3 in SM sm . To proceed systematically, setting aside for the time being the question of which of the two defects will have a DW behind it, let us separately consider DW formations by PD and ED. Considering a PD, a DW is first formed in $T_{1}$, as $n$ just exceeds $n_{cL}^{p}$, a threshold for DW formation (see below) with the DW position $x_{w}$ being at $x=0$. As $n$ exceeds $n_{cL}^{p}$, the size of this pile or the DW increases by shifting its position $x_{w}$ (see below for a formal definition) along the ring, moving from $T_{1}$ first to $T_{2}$ and then to $T_{3}$, finally reaching $x=1$, at which point the DW ends at a threshold $n_{cU}^{p}$ of $n$ (see below), bringing the model to its HD-HD-HD phase. Now assume a DW at $x_{w}$ in $T_{1}$, i.e., $0<x_{w}<\epsilon_{1}$, connecting a high density segment with density $\rho_{\text{HD}}=1/(1+p)$ between 00 and $x_{w}$ and a low density segment with density $\rho_{\text{LD}}=p/(1+p)$ between $x_{w}$ and $\epsilon_{1}$ lebo ; niladri-tasep . Then using PNC
@@
Clearly, these solutions are physically acceptable, so long as $14p/[q(1+p)^{2}]<$ 1. For a DW in $T_{1}$, $\rho_{2}=\rho_{2-}$. Then applying PNC
@@
a lower threshold on $n$, such that as $n$ exceeds $n_{cL}^{p}$ a DW due to PD forms in $T_{1}$ with the disappearance of the LD-LD-LD phase. As $n$ rises, the LDW shifts first to $T_{2}$ and then to $T_{3}$. The corresponding locations can be found by using PNC in a way analogous to the above. In particular, an LDW in $T_{3}$ has a location
@@
An LDW can be formed in $T_{2}$ also; see SM sm for the corresponding LDW position. At $x_{w}=1$, the LDW ends, giving the boundary between the HD-HD-DW and HD-HD-HD phases. We find
@@
an upper threshold on $n$, such that as $n$ exceeds $n_{cU}^{p}$, HD-HD-HD phase reappears. See SM sm for a similar expression of $x_{w}$ in $T_{2}$. In fact, an LDW in $T_{2}$ assumes a staircase-like shape. See Fig. 3 (left) for a plot of an LDW in $T_{2}$.
Refer to captionRefer to caption
@@
A DW can however also form behind ED, once a rising $n$ exceeds a different lower threshold, denoted $n_{cL}^{q}$ below. At $n=n_{cL}^{q}$, $T_{2}$ reaches its MC phase with $\rho_{2}=1/2$ in the bulk, and $\rho_{1}=\rho_{3}<1/2$. As $n$ is further increased beyond $n_{cL}^{q}$, a DW is formed in the system, just behind the bottleneck. In the ED dominated regime considered here, $T_{2}$ is the (extended) bottleneck.This means a DW is first formed at $x=\epsilon_{1}+\epsilon_{2}$ in $T_{3}$ as soon as $n$ reaches $n_{cL}^{q}$. As $n$ rises further, the DW starts moving along $T_{3}$, crossing over to $T_{1}$ at $x=0$, and then finally reaching $x=\epsilon_{1}$ at an upper density threshold $n_{cU}^{q}$, when the DW ends. Any further increase in $n$ will push the system to the HD-HD-HD phase. Thus, the DW never enters $T_{2}$, which remains in its MC phase. This is a fundamental difference with the LDW formed due to PD.
@@
giving the densities of the high density and low density parts of the DW, which meet at $x_{w}$. Using the logic outlined above for a PD together with (7) and PNC, we find with $x_{w}=\epsilon_{1}+\epsilon_{2}$
@@
defining an upper threshold on $n$, such that for $n>n_{cU}^{q}$, HD-HD-HD phase ensues. Our MFT and MCS results on LDW in $T_{1}$ is shown in Fig. 3 (right). Following the logic outlined above the LDW position in $T_{1}$ or $T_{3}$ due to ED can be found straightforwardly sm .
A minimum current principle determines which of the two possible routes to DW formations is actually realised for a given $(p,q)$. The stationary current $J_{\text{PD}}$ corresponding to a domain wall solution due to PD is $J_{\text{PD}}={p}/{(1+p)^{2}}$. Similarly, the stationary current $J_{\text{ED}}$ corresponding corresponding to a domain wall due to ED is $J_{\text{ED}}=\frac{q}{4}$. Minimum current principle stipulates that when $J_{\text{PD}}<(>)J_{\text{ED}}$, the nonuniform steady states are controlled by the PD (ED), which is consistent with our MCS results. When $J_{\text{PD}}=J_{\text{ED}}$, the two defects compete, as we illustrate below, a new kind of states - a pair of delocalised domain walls (DDW) emerges as long as $n$ exceeds both $n_{cL}^{q},\,n_{cL}^{p}$, but less than $n_{cU}^{q},\,n_{cU}^{p}$. Following the logic of LDW formation in PD and ED dominated regimes, we then expect one DW, say at $x=x_{w}^{\text{ED}}$, due to ED, and another, $x=x_{w}^{\text{PD}}$, due to PD. Thus, a complete description of the stationary density profiles require enumeration of both $x_{w}^{\text{PD}}$ and $x_{w}^{\text{ED}}$. However, with just one condition, viz., the PNC at hand, both $x_{w}^{\text{PD}}$ and $x_{w}^{\text{ED}}$ cannot be uniquely determined. Rather a linear relation between the two is obtained. This means any pair of ($x_{w}^{\text{PD}}$, $x_{w}^{\text{ED}}$) satisfying PNC is a valid solution. Since a DW due to ED must be confined to $T_{1}$ and $T_{3}$ only (see above), other DW due to PD must also be confined to the same (even though an LDW due to an isolated PD with $J_{\text{PD}}<J_{\text{ED}}$ can also be in $T_{2}$). Thus, each of ($x_{w}^{\text{PD}}$, $x_{w}^{\text{ED}}$) must be confined to $T_{1}$ or $T_{3}$ only. The inherent stochasticity of the dynamics implies all such ($x_{w}^{\text{PD}}$, $x_{w}^{\text{ED}}$) pairs satisfying PNC are visited over time. This further means long-time averages of the densities are inclined straight lines. As a concrete example, assume $n_{cL}^{p},\,n_{cL}^{q}<n<1/2$ together with $J_{\text{ED}}=J_{\text{PD}}$. Both the DWs must have the same height $(1-p)/(1+p)$, with $\rho_{\text{HD}}=1/(1+p),\,\rho_{\text{LD}}=p/(1+p),\,\rho_{2}=1/2$. Applying PNC,
@@
giving a linear relation connecting $x_{w}^{\text{PD}}$ and $x_{w}^{\text{ED}}$, and not each of the positions separately. As a result, a pair of DDWs is observed. MFT cannot predict the profiles of the DDWs, as it neglects fluctuations. However, we can employ arguments based on symmetry to construct the DDW profiles. Focussing on the case $\epsilon_{1}=1/3=\epsilon_{2}$ and exploiting the statistical equivalence of the configurations of the long time averaged envelop of the DDWs in $T_{1}$ and $T_{3}$, we hypothetically replace each of them by an LDW of height $(1-p)/(1+p)$, connecting $\rho_{\text{HD}}=1/(1+p)$ and $\rho_{\text{LD}}=p/(1+p)$. Such (hypothetical) replacements evidently satisfy the PNC. Symmetry of the problem dictates that if $x_{0}$ is the position of the LDW in $T_{1}$, the corresponding LDW in $T_{3}$ must be located at $x_{0}+2/3$. Application of PNC then gives
@@
Clearly, if $n=1/2$, $x_{0}=1/6$, which means the midpoint of the LDW is at the midpoint of $T_{1}$ or $T_{3}$, corresponding to DDWs covering entire $T_{1}$ and $T_{3}$. If $n<1/2$, $x_{0}<1/6$, whereas $n>1/2$, $x_{0}>1/6$, both of which correspond to DDWs partially covering $T_{1}$ and $T_{3}$. With the knowledge of $x_{0}$, it is now possible to obtain the DDW envelop. For instance with $n<1/2$, there is an LD segment in the DDWs in both $T_{1}$ and $T_{3}$, which are of length $1/3-2x_{0}$. In other words, the DDW in $T_{1}$ ($T_{3}$) wanders a distance $2x_{0}$, starting from $x=0$ ($x=\epsilon_{1}+\epsilon_{2}$). Then joining the densities at $x=0,2x_{0}$ ($x=\epsilon_{1}+\epsilon_{2},\epsilon_{1}+\epsilon_{2}+2x_{0}$) gives the profiles of the DDW in $T_{1}$ ($T_{2}$). Our MCS results on DDWs in $T_{1},\,T_{3}$ and the corresponding kymographs are shown in Fig. 4. The analytically obtained DDW profiles are also shown. The kymographs clearly show the synchronised nature of the DDW movements fully or partially covering $T_{1}$ and $T_{3}$. See sm for a related movie (Movie 4) that visually presents this picture, showing a pair of DDWs, partial or full as controlled by $n$.
Refer to captionRefer to caption
Refer to captionRefer to caption
@@
Figure 4: Plot of a pair of DDWs in $T_{1}$ and $T_{3}$ with (top left) $n=0.5,\,p=0.2,\,q=0.5933,\,L=9000,\,\epsilon_{1}=\epsilon_{2}=1/3$, (top right) $n=0.35,\,p=0.2,\,q=0.5933,\,L=9000,\,\epsilon_{1}=\epsilon_{2}=1/3$, and corresponding kymographs (bottom left) and (bottom right).The analytically obtained DDW profiles (blue lines) and MCS results (red points) are shown (see text).

### v3

@@ Abstract
Totally asymmetric simple exclusion process (TASEP) sets the paradigm for one-dimensional driven single file motion. We study a periodic TASEP with two “road blocks” or defects of different kinds, one point and another extended, across which particle flows are inhibited. We show how the interplay between particle number conservation and competition between the defects lead to inhomogeneous steady states with localised domain walls (LDW). The LDW locations jump discontinuously, indicating a discontinuous transition between these LDW states, as the system passes from being controlled by one defect to the other. When the defects are “competing”, instead of an LDW a pair of delocalised domain walls appear, none of which can penetrate the extended defect. A minimum current principle can be used to identify the dominant defect that controls the domain wall formations. Our results should be important in diverse systems, ranging from protein synthesis by ribosomes in biological cells to urban traffic networks.
I Introduction
@@
Single file motion implies particle motion along quasi one-dimensional (1D) narrow channels where the particles cannot cross each other due to hardcore repulsion. This was originally introduced by Hodgkin and Keynes hodg to describe ion transport in biological channels. Unidirectional or driven single file motion have received increasing attention in the recent past for their wide-ranging interdisciplinary applications. We study the role of bottlenecks and their interplay and competition in controlling the stationary states of driven single file motion, a question of potential relevance in wide-ranging systems, e.g., traffic flow  ped1 ; ped2 ; ped3 , colloidal particles flowing through geometric constraints coll1 ; coll2 ; coll3 and ribosome translation along messenger RNA (mRNA) loops mrna1 .
Totally asymmetric simple exclusion process (TASEP), a paradigmatic example of 1D driven single file motion, consists of 1D lattice with $L$ sites, with particles hopping unidirectionally, subject to exclusion. Originally proposed to describe protein synthesis by ribosome translocation along messenger RNA strands in eukaryotic cells mac , it was re-invented as a archetype for nonequilibrium phase transition in 1D open systems krug ; tom . TASEP has been adopted to gain generic understanding about protein synthesis in cells mrna1 ; tom-prl ; beate1 . In particular, Refs. tom-prl ; beate1 ; krug-new explored the sensitive dependence of the particle (“ribosome”) currents in an open TASEP with isolated or clusters of bottlenecks (“rare codons”), theoretical results that resonate with relevant experimental findings bao-natcomm ; see also Ref. mrna5 . In a surprising cross-field connection, congestion in vehicular or pedestrian movements by road blocks - increasingly important in urban life, has also been addressed using TASEP-like cellular automata models with bottlenecks comp . While the effects of defects on the stationary currents in an open TASEP has already been studied, both in the context of protein synthesis in cells  tom-prl ; beate1 ; krug-new and also from pure nonequilibrium physics standpoint open1 ; open2 ; open3 ; atri1 ; sm-ab-tasep , the question of identifying the dominant defect and its interplay in a closed many-defect transport system remains open.
@@
In this Article, we study a particle number conserving TASEP in a ring with one point and one extended defects as a conceptual model for competition among road blocks in a closed driven single file motion. We elucidate how the stationary density profiles of the TASEP emerge as an interplay between the dominant defect and number conservation. Our principal results are (i) the defects lead to domain walls (DW) in the stationary density profiles connecting segments lower and higher than 1/2,when the system is neither nearly empty or nearly full, but rather has a moderate density. This defines a DW phase. In a striking display of defect competition, the DW location jumps discontinuously as the system undergoes a hitherto unstudied first order transition between point defect (PD) and extended defect (ED) controlled states. (ii) In the DW phase, generally there is a single localised DW (LDW) for both PD or ED. Nonetheless, the LDW profiles for a PD and ED are fundamentally different from each other. (iii) In the event, the two defects are equi-dominant (in the sense explained below), instead of a single LDW, a pair of delocalised domain walls (DDW) emerges, forming outside ED. The latter is in its maximal current (MC) phase. (iv) For a nearly empty or filled system, the density everywhere is less or more than 1/2, reminiscent of the low density (LD) or high density (HD) phases of a TASEP with open boundary conditions krug ; tom ; rev1 . Our results on DWs can be reconciled within a minimum current principle.
Our work should be useful for understanding the phenomenologies of ribosome translocations along closed mRNA loops (circular translation of polysomal mRNA) with clusters of slow codons along which ribosome translocations are inhibited mrna1 ; mrna11 ; mrna2 ; mrna3 ; mrna4 ; mrna5 and traffic jams or congestion in urban transport networks with multiple road blocks ped2 ; ped3 . From a theoretical standpoint, our results complement existing research on the effects of a slow bond on 1D driven systems; see Refs. soh ; soh1 .
II Model
@@
The phases in the model and the associated transitions are controlled by $p,q,n$. The mean-field phase diagram with $\epsilon_{1}=\epsilon_{2}=1/3$ is shown in Fig. 2, giving the phases in the three segments - $T_{1}$ between $i=1$ to $\epsilon_{1}L$, $T_{2}$ between $\epsilon_{1}L$ to $(\epsilon_{1}+\epsilon_{2})L$ and $T_{3}$ between $(\epsilon_{1}+\epsilon_{2})L$ to $L$, connected serially. See Fig. 3 (left) with $p=0.20,q=0.65$ and Fig. 3 (right) with $p=0.20,q=0.55$ for the associated fundamental diagrams, giving the current-density relations, in the PD and ED dominated regions of the phase space, obtained from mean-field theory (MFT) and Monte-Carlo simulations (MCS).
Figure 2: Mean-field phase diagram of the model in the $p-q-n$ space showing the phases in the model. HD and DW (with one LDW) phases are marked. LD phase that lies below the DW phase in the figure is not visible. Points on the gray curved plane lying as a divider of the DW phase have $J_{\text{PD}}=J_{\text{ED}}$ satisfied, where $J_{\text{PD}},\,J_{\text{ED}}$ are the steady state currents in the PD and ED dominated situations respectively (see text below) and have a pair of DDWs fully or partially covering $T_{1}$ and $T_{3}$. In the DW phase region, on the left of this curved surface (i.e., with $J_{\text{PD}}<J_{\text{ED}}$), there is an LDW due to the point defect in either $T_{1},T_{2}$ or $T_{3}$; on the right of this plane (i.e., with $J_{\text{ED}}<J_{\text{PD}}$), one has an LDW due to the extended defect in $T_{1}$ or $T_{3}$ and correspondingly an MC phase in $T_{2}$. In the HD (LD) phase region, one has HD-HD-HD (LD-LD-LD) phase in $T_{1},T_{2},T_{3}$ (see text).
Figure 3: Fundamental diagrams ($\epsilon_{1}=\epsilon_{2}=1/3$) the phase space regions dominated by the (left) point defect ($p=0.2,q=0.65$), with visible mismatch between MFT and MCS results and some finite size effects, (right) extended defect ($p=0.2,q=0.55$), with good MFT-MCS agreement and no observable finite size effects (see text).
@@
III Domain Walls
In this section, we focus on the DWs. We derive the corresponding results on the DWs within MFT blythe , supplemented by extensive MCS studies, which provide quantitative basis to the movies in SM. The density profiles for the phases without any DWs, i.e., in LD-LD-LD and HD-HD-HD phases, can be obtained by using particle number and current conservations; see Appendix A. In addition, the absence of MC-MC-MC phase is argued in Appendix A.
@@ III Domain Walls
The emergence of DWs, as revealed in our MCS results and observed in the movies sm can happen in two distinct ways. Physically, from the LD-LD-LD phase upon addition of particles, $n$ rises and it eventually reaches a lower threshold, a macroscopically nonuniform steady state in the shape of a DW is formed. A DW, essentially a pile up of particles due to a bottleneck or a defect, should form behind the defect, point or extended, as the case may be, and then start to grow as $n$ rises further. While this remains true independent of the defect, our MCS results reveal a striking aspect - a DW that forms behind a PD can be anywhere in the TASEP ring covering all of $T_{1},\,T_{2},\,T_{3}$, being controlled by $n$, whereas a DW, which forms behind an ED, never enters $T_{2}$! This leads to complex density profiles when the two defects are “competing”; see below and also Movie 3 in SM sm . To proceed systematically, setting aside for the time being the question of which of the two defects will have a DW behind it, let us separately consider DW formations by PD and ED. Considering a PD, a DW is first formed in $T_{1}$, as $n$ just exceeds $n_{cL}^{p}$, a threshold for DW formation (see below) with the DW position $x_{w}$ being at $x=0$. As $n$ exceeds $n_{cL}^{p}$, the size of this pile or the DW increases by shifting its position $x_{w}$ (see below for a formal definition) along the ring, moving from $T_{1}$ first to $T_{2}$ and then to $T_{3}$, finally reaching $x=1$, at which point the DW ends at a threshold $n_{cU}^{p}$ of $n$ (see below), bringing the model to its HD-HD-HD phase. Now assume a DW at $x_{w}$ in $T_{1}$, i.e., $0<x_{w}<\epsilon_{1}$, connecting a high density segment with density $\rho_{\text{HD}}=1/(1+p)$ between 00 and $x_{w}$ and a low density segment with density $\rho_{\text{LD}}=p/(1+p)$ between $x_{w}$ and $\epsilon_{1}$ lebo ; niladri-tasep . Then using number conservation
@@
These solutions are physically acceptable for $14p/[q(1+p)^{2}]<$ 1. For a DW in $T_{1}$, $\rho_{2}=\rho_{2-}$. Number conservation then gives
@@
a lower threshold on $n$, such that as $n$ exceeds $n_{cL}^{p}$ a DW due to PD forms in $T_{1}$ with the disappearance of the LD-LD-LD phase. As $n$ rises, the LDW shifts first to $T_{2}$ and then to $T_{3}$. The corresponding locations can be found by using the logic outlined above. In particular, for an LDW in $T_{3}$
@@
An LDW can be formed in $T_{2}$ also; see Appendix B for the corresponding LDW position. In fact, an LDW in $T_{2}$ assumes a staircase-like shape. See Fig. 4 (left) for a plot of an LDW in $T_{2}$. At $x_{w}=1$, the LDW ends, giving the boundary between the HD-HD-DW and HD-HD-HD phases. We find
@@
an upper threshold on $n$, such that as $n$ exceeds $n_{cU}^{p}$, HD-HD-HD phase appears.
@@
A DW can however form behind ED, once $n$ exceeds a different lower threshold $n_{cL}^{q}$. At $n=n_{cL}^{q}$, $T_{2}$ reaches its MC phase with $\rho_{2}=1/2$ in the bulk, and $\rho_{1}=\rho_{3}<1/2$. As $n$ is further increased beyond $n_{cL}^{q}$, a DW is formed in the system, just behind the bottleneck. In the ED dominated regime considered here, $T_{2}$ is the (extended) bottleneck.This means a DW is first formed at $x=\epsilon_{1}+\epsilon_{2}$ in $T_{3}$ as soon as $n$ reaches $n_{cL}^{q}$. As $n$ rises further, the DW starts moving along $T_{3}$, crossing over to $T_{1}$ at $x=0$, and then finally reaching $x=\epsilon_{1}$ at an upper density threshold $n_{cU}^{q}$, when the DW ends. Any further increase in $n$ will push the system to the HD-HD-HD phase. Thus, the DW never enters $T_{2}$, which remains in its MC phase. This is a fundamental difference with the LDW formed due to PD.
@@
giving the densities of the high density and low density parts of the DW, which meet at $x_{w}$. Using the logic outlined above for a PD together with (7) and particle number conservation, we find with $x_{w}=\epsilon_{1}+\epsilon_{2}$
@@
defining an upper threshold on $n$, such that for $n>n_{cU}^{q}$, HD-HD-HD phase ensues. Our MFT and MCS results on LDW in $T_{1}$ is shown in Fig. 4 (right). Following the logic outlined above the LDW position in $T_{1}$ or $T_{3}$ due to ED can be found straightforwardly; see Appendix C.
A minimum current principle determines which of the two possible routes to DW formations is actually realised for a given $(p,q)$. The stationary current $J_{\text{PD}}$ corresponding to a domain wall solution due to PD is $J_{\text{PD}}={p}/{(1+p)^{2}}$. Similarly, the stationary current $J_{\text{ED}}$ corresponding to a domain wall due to ED is $J_{\text{ED}}=\frac{q}{4}$. Minimum current principle stipulates that when $J_{\text{PD}}<(>)J_{\text{ED}}$, the nonuniform steady states are controlled by the PD (ED), which is consistent with our MCS results. When $J_{\text{PD}}=J_{\text{ED}}$, the two defects compete, as we illustrate below, a new kind of states - a pair of delocalised domain walls (DDW) emerges for $n_{cL}^{q},\,n_{cL}^{p}<n<n_{cU}^{q},\,n_{cU}^{p}$. Following the logic of LDW formation in PD and ED dominated regimes, we then expect one DW, say at $x=x_{w}^{\text{ED}}$, due to ED, and another, $x=x_{w}^{\text{PD}}$, due to PD. Thus, a complete description of the stationary density profiles require enumeration of both $x_{w}^{\text{PD}}$ and $x_{w}^{\text{ED}}$. However, with just one condition, viz., particle number conservation, both $x_{w}^{\text{PD}}$ and $x_{w}^{\text{ED}}$ cannot be uniquely determined. Rather a linear relation between the two is obtained. This means any pair of ($x_{w}^{\text{PD}}$, $x_{w}^{\text{ED}}$) satisfying particle number conservation is a valid solution. Since a DW due to ED must be confined to $T_{1}$ and $T_{3}$ only (see above), other DW due to PD must also be confined to the same (even though an LDW due to an isolated PD with $J_{\text{PD}}<J_{\text{ED}}$ can also be in $T_{2}$). Thus, each of ($x_{w}^{\text{PD}}$, $x_{w}^{\text{ED}}$) must be confined to $T_{1}$ or $T_{3}$ only. The inherent stochasticity of the dynamics implies all such ($x_{w}^{\text{PD}}$, $x_{w}^{\text{ED}}$) pairs satisfying particle number conservation are visited over time. This further means long-time averages of the densities are inclined straight lines. Both the DWs must have the same height $(1-p)/(1+p)$, with $\rho_{\text{HD}}=1/(1+p),\,\rho_{\text{LD}}=p/(1+p),\,\rho_{2}=1/2$. Particle number conservation gives
@@
giving a linear relation connecting $x_{w}^{\text{PD}}$ and $x_{w}^{\text{ED}}$, and not each of the positions separately. As a result, a pair of DDWs is observed. MFT cannot predict the profiles of the DDWs, as it neglects fluctuations. However, we can employ arguments based on symmetry to construct the DDW profiles. With $\epsilon_{1}=1/3=\epsilon_{2}$ and exploiting the statistical equivalence of the configurations of the long time averaged envelope of the DDWs in $T_{1}$ and $T_{3}$, we hypothetically replace each of them by an LDW of height $(1-p)/(1+p)$, connecting $\rho_{\text{HD}}=1/(1+p)$ and $\rho_{\text{LD}}=p/(1+p)$. Such replacements evidently satisfy particle number conservation. Symmetry of the problem dictates that if $x_{0}$ is the position of the LDW in $T_{1}$, the corresponding LDW in $T_{3}$ must be located at $x_{0}+2/3$. Application of particle number conservation gives
@@
Clearly, if $n=1/2$, $x_{0}=1/6$, which means the midpoint of the LDW is at the midpoint of $T_{1}$ or $T_{3}$, corresponding to DDWs covering entire $T_{1}$ and $T_{3}$. If $n<1/2$, $x_{0}<1/6$, whereas $n>1/2$, $x_{0}>1/6$, both of which correspond to DDWs partially covering $T_{1}$ and $T_{3}$. With the knowledge of $x_{0}$, it is now possible to obtain the DDW envelope. For instance with $n<1/2$, there is an LD segment in the DDWs in both $T_{1}$ and $T_{3}$, which are of length $1/3-2x_{0}$. In other words, the DDW in $T_{1}$ ($T_{3}$) wanders a distance $2x_{0}$, starting from $x=0$ ($x=\epsilon_{1}+\epsilon_{2}$). Then joining the densities at $x=0,2x_{0}$ ($x=\epsilon_{1}+\epsilon_{2},\epsilon_{1}+\epsilon_{2}+2x_{0}$) gives the profiles of the DDW in $T_{1}$ ($T_{2}$). Our MCS results on DDWs in $T_{1},\,T_{3}$ and the corresponding kymographs are shown in Fig. 5 (top, left) and (top, right) and Fig. 5 (bottom, left) and (bottom, right), respectively. The analytically obtained DDW profiles are also shown. The kymographs in Fig. 5 (bottom, left) and Fig. 5 (bottom, right) clearly show the synchronised nature of the DDW movements fully or partially covering $T_{1}$ and $T_{3}$. See sm for a related movie (Movie 4) that visually presents this picture, showing a pair of DDWs, partial or full as controlled by $n$.
@@
Figure 5: Plots of a pair of DDWs in $T_{1}$ and $T_{3}$ with parameters (top, left) $n=0.5,\,p=0.2,\,q=0.5933,\,L=9000,\,\epsilon_{1}=\epsilon_{2}=1/3$, (top, right) $n=0.35,\,p=0.2,\,q=0.5933,\,L=9000,\,\epsilon_{1}=\epsilon_{2}=1/3$, and corresponding kymographs (bottom, left) and (bottom, right).The analytically obtained DDW profiles (blue lines) and MCS results (red points) are shown (see text).